# Optical Physiological Signals

### v1

@@ 4.1 Pre-training
Datasets. We curate three large public datasets to pre-train PaPaGei (as shown in Table 1). The first dataset is VitalDB (Lee et al., 2022), which includes PPG signals collected during surgery from the finger using a patient monitor ($f$=500Hz). The second dataset is the MIMIC-III waveform database matched subset (Johnson et al., 2016), where finger-tip PPG data is collected from an ICU monitor ($f$= 125Hz). The third dataset comes from the Multi-Ethnic Study of Atherosclerosis (MESA) sleep sub-study (Zhang et al., 2018; Chen et al., 2015), which provides PPG data obtained through finger-tip polysomnography ($f$= 256Hz). In total, we have 13.5K subjects with 20M segments.
Table 1: Pre-training datasets.
Dataset #Subjects #Segments Hours
@@
Pre-processing. To curate and harmonize the PPG signals across all datasets, we perform the following steps based on established practices: (1) Apply a 4th-order Chebyshev bandpass filter with low and high pass cut-offs set at 0.5Hz and 12Hz, respectively (Lapitan et al., 2024; Liang et al., 2018c); (2) Segment the signal into 10-second windows (Orphanidou, 2018; Koteska et al., 2022); (3) Detect flatline segments and remove any segment where more than 25% of the data is flat (BioBSS Documentation, 2023); (4) Normalize the segments using Z-score (Abbaspourazad et al., 2023; Temko, 2017; Zhou et al., 2017); and (5) Resample the segments to 125Hz (the lowest sampling rate of our pre-training datasets, MIMIC-III).
Table 2: The task evaluation benchmark of PaPaGei. Datasets in grey are unseen during training (out-of-domain). The rest are used for pre-training (held-out test-sets and labels). Task types in brackets are B: binary, R: regression, M-#classes: multiclass classification.
#ID Dataset Task #Subj.(#Samp.)
@@
Implementation. Recent work in PPG modeling has extensively used small-scale CNN architectures (Abbaspourazad et al., 2023; Ding et al., 2024). Building on that, we implement a ResNet-style (He et al., 2016) CNN encoder with 18 convolutional blocks, starting with a filter size of 32, which doubles every 4 blocks. The projection layer is a single FC layer, generating a 512-d embedding. In the PaPaGei-S variant, the expert block ($M_{1}$ $M_{2}$) uses three parallel FCNNs, each with two FC layers, resulting in a 128-d embedding. For augmentations, PaPaGei-P uses cropping (0.50), negation (0.20), flipping (0.20), and scaling (0.40). PaPaGei-S uses cropping (0.25) and Gaussian noise (0.25). PaPaGei-S avoids augmentations that alter PPG’s morphology. We set $\alpha=0.6$ and train on eight V100 GPUs for 15,000 steps (lr$=10^{-4}$), with PaPaGei-P and PaPaGei-S having 5M and 5.7M parameters, respectively (we study scaling in Section 5.2).
@@ 4.2 Downstream Tasks
To evaluate the effectiveness of PaPaGei, we benchmark it against a diverse set of datasets, tasks, and baselines, chosen for their large size and clinical relevance (where applicable) [^2] https://peterhcharlton.github.io/post/ppg_datasets/. A brief description of the tasks with their corresponding #ID is provided in Table 2, with further details in Appendix B.
@@
Identifying patient risk factors is crucial for hospitals to allocate resources effectively. To address this, we evaluate several indicators, including ICU admission (T1), type of operation (T2), mortality (T3), and smoking status (T4). For sleep apnea diagnosis, the American Academy of Sleep Medicine (AASM) recommends using the Apnea/Hypopnea Index (AHI) with at least 3% or 4% oxygen desaturation as a key metric (Ruehland et al., 2009). Thus, we predict AHI at 3% and 4% desaturation thresholds (T5 & T6) and classify sleep-disordered breathing (T15). For pregnancy outcomes, changes in gestational age and pregnancy stage are linked to risks like hypertensive disorders and small-for-gestational-age delivery (Bouariu et al., 2022; Wu et al., 2020; Crump et al., 2023), enabling us to classify pregnancy stage (T7) and predict gestational age (T8). In cardiovascular health, we estimate systolic (T9 & T11) and diastolic (T10 & T12) blood pressure (BP) using two datasets. While PPG-BP (T11 & T12) provides high-frequency, short PPG signals, the VV dataset helps explore skin tone’s influence on BP estimation. We also assess hypertension classification (T14), average seated heart rate (T13), and continuous heart rate during activities (T19), along with activity classification (T20). In the emotion domain, we classify PPG signals into mood disturbance levels (T16), valence (T17), and arousal (T18).
@@ 4.3 Baselines
We benchmark PaPaGei’s performance against competitive baselines. As open-source foundation models designed for physiological signals, PaPaGei is compared to recent time-series foundation models: Chronos (Ansari et al., 2024) and MOMENT (Goswami et al., 2024). To evaluate the merits of our SSL framework, we also compare PaPaGei with common SSL methods (trained from scratch) such as SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), and TF-C (Zhang et al., 2022). In addition, to assess model generalizability on PPG signals, we compare against REGLE, a model pre-trained on UK Biobank’s PPG signals (Yun et al., 2024). As a simple baseline, we employ a random forest trained on statistical features extracted from the PPG signal, including mean, median, maximum, minimum, and the 25th, 50th, and 75th percentiles (”Stat. Features”). This task-specific approach serves as a benchmark for comparison with more advanced techniques.
@@ 4.4 Linear Evaluation
Initially, we split the in-domain and out-of-domain datasets into training, validation, and test sets at 80/10/10 and 60/20/20 ratios. The splitting is performed at the subject level ensuring no overlap between individuals across the sets. The models are evaluated by extracting feature representations from resampled data (125Hz) and applying linear probing for each task. For binary classification tasks, we employ a logistic regression model, with performance measured by the AUROC score. For regression tasks, ridge regression is used, and performance is evaluated based on the mean absolute error (MAE). Multi-class classification tasks are trained using a random forest model, with accuracy as the evaluation metric. To ensure robustness, we compute 95% confidence intervals (CIs) through bootstrapping (500 sampling runs with replacement). More details are provided in the Appendix §A.
@@ 5.1 Overall Performance
Table 3: Downstream comparison against pre-trained models. The model size is denoted next to their names. 95% CIs are reported in square brackets and the best value is bolded.
REGLE (0.07M) Chronos (200M) Moment (385M) PaPaGei-P (5M) PaPaGei-S (5.7M)
@@
Average 12.09 $\pm$ 3.83 11.60 $\pm$ 3.60 10.43 $\pm$ 3.46 10.92 $\pm$ 4.25 10.12 $\pm$ 3.47
In general, from Figure 3, we observe that PaPaGei is more accurate across many tasks indicated by the larger AUROC area and smaller MAE area. Table 3 presents a more detailed comparison between PaPaGei and other pre-trained models. For classification tasks, PaPaGei-S achieves the highest average AUROC of 0.67, outperforming other models across several tasks, particularly in ICU Admission (0.79), Hypertension (0.77), and Sleep Disordered Breathing (0.70). In regression tasks, PaPaGei-S again demonstrates strong performance, achieving the lowest average MAE (10.12), particularly in tasks related to Apnea/Hypopnea Index and BP measurements. REGLE, a small model trained on a large PPG dataset, generally underperforms compared to other models, suggesting its compact size may limit learning complex patterns. Chronos, a much larger model, shows competitive performance, particularly in Pregnancy Stage and Gestation Age prediction. Moment, the largest model, shows strong performance across many tasks, often matching or slightly outperforming Chronos. Notably, while PaPaGei-P shows competitive results, PaPaGei-S consistently outperforms it, suggesting the signal morphology objectives offer advantages in predictive capability.
Table 4 presents a comparison against three SSL methods and a baseline model trained on statistical features. In classification tasks, PaPaGei-S again shows the highest average AUROC, outperforming all others. SimCLR, BYOL, and TF-C generally outperform the statistical feature baseline but fall short of PaPaGei-S’s performance. TF-C shows competitive results in some tasks, achieving the highest AUROC for Mood Disturbance and Valence.
@@
For regression tasks, PaPaGei-S again achieves the lowest average MAE. SimCLR, BYOL, and TF-C show mixed results, as each excels in different tasks. SimCLR comes second in estimating Avg HR, while BYOL does so in Systolic BP (VV). The statistical feature baseline generally underperforms compared to the advanced methods across most tasks. PaPaGei-P, while not consistently outperforming PaPaGei-S, shows strong results that are often competitive with or better than other contrastive learning methods. Overall, both PaPaGei variants offer robust performance across a wide range of tasks.
Table 4: Downstream comparison against baseline and other SSL methods. The model size is denoted next to their names. 95% CIs are reported in square brackets and the best value is bolded.
Stat. Features SimCLR (5M) BYOL (12M) TF-C (9.7M) PaPaGei-P (5M) PaPaGei-S (5.7M)
@@ 5.2 Ablation Studies
PaPaGei-S component ablation. We assess the impact of PaPaGei-S components. Figure 5(a) & 5(b) show the full model consistently outperforms individual components in both mean and median metrics. On average, sVRI (0.64, 10.36) outperforms the combinations of sVRI + SQI (0.62, 10.81) and sVRI + IPA (0.64, 10.73), with SQI contributing the least to overall performance.
@@
Figure 5: PaPaGei-S component ablation study (a, b) and scaling analysis (c). (Left) The boxplot shows the performance of PaPaGei-S components across all tasks. Despite the wide variation, the full model outperforms individual components. (Right) Heatmap ranks of PaPaGei-S models with 5M, 35M, and 139M parameters (rank 1 denotes the best performance). Detailed results in Table 13.
Model size and scaling analysis. We evaluated whether larger model sizes improve performance by additionally training PaPaGei-S-35M and PaPaGei-S-139M with 35M and 139M parameters, respectively. Both models have the same number of layers, but the 35M model uses a 32 starting filter size, while the 139M model uses 64. As shown in 5(c), the smallest model with 5M parameters performs best in all tasks except one, indicating that smaller models are better suited for PPG data.
Figure 6: Ablation on pre-training datasets. Average performance across tasks for Binary classification (left) and regression (right) for models trained on different datasets: V (VitalDB), M (MESA), and M-III (MIMIC-III). The mean value is displayed above the plots.
Notably, the 139M model outperforms the 35M but still falls short of the 5M, suggesting that wider models may offer improved performance. Furthermore, the added complexity of the 139M model is advantageous for capturing decision boundaries in classification compared to the 35M model, which could be attributed to the classification-like objectives of contrastive learning. However, it struggles with regression tasks, which demand more precise outputs.
Pre-training data ablation. We evaluated PaPaGei-S using different combinations of pre-training data to determine if increasing the data volume enhances performance. Figure 6 demonstrates that performance on downstream tasks consistently improves as we incorporate additional upstream data, with optimal results achieved when utilizing all three datasets. Interestingly, in single datasets, MESA performs the best, despite having the fewest participants but the highest number of segments, echoing similar results in the language model domain (Dubey et al., 2024).
@@ 5.3 Case Studies
Figure 7: Pair-wise inter-participant embedding distances in sleep-disordered breathing.
To evaluate the downstream performance of PaPaGei-S, we analyze the embedding distances among users within the sleep-disordered breathing dataset. We also investigate the model’s performance in predicting the Apnea-Hypopnea Index (AHI $>$ 3%) and average Heart Rate. We analyze BP performance across different skin tones to evaluate robustness to different populations.
@@
Inter-participant embeddings. Figure 7 shows the distribution of pair-wise embedding distances across participants (Kiyasseh et al., 2021). SimCLR and BYOL exhibit sharper peaks at lower distances, indicating that participants are more closely clustered within the embedding space. This could be interpreted as a mild form of mode collapse, where the model does not fully capture the individual differences between participants. TF-C demonstrates a more balanced distribution, with both large and small peaks, suggesting it captures both similarities and some variation between participants. In contrast, PaPaGei-S provides the widest dispersion of embeddings, highlighting its ability to capture a broader range of features that may be valuable for distinguishing between participants’ medical conditions.
((a)) AHI $>$ 3% predictions
((b)) Avg. HR predictions
Figure 8: Regression plots with histograms for (a) Apnea/Hypopnea $>$ 3% and (b) HR predictions. PaPaGei consistently matches the true label (scalar) distribution.
Figure 9: Skin tone analysis for Blood Pressure estimation (VV dataset). Figure 22 provides a breakdown across all skin tones.
Regression predictions. In Figure 8, we observe that all models demonstrate positive correlations between true and predicted AHI. Compared to Chronos and SimCLR, PaPaGei-S shows superior regression performance (r=0.54 vs. r=0.44 and r=0.41). This suggests that PaPaGei-S provides a more accurate prediction of AHI $>$ 3% compared to Chronos and SimCLR, as it aligns more closely with the true AHI values. Moreover, PaPaGei-P demonstrates a higher correlation (0.90 vs. 0.77 and 0.83) for HR prediction. Examining the PaPaGei-P histograms reveals that the model effectively captures both tails of the distribution, reducing errors in the high and low heart rate ranges. In contrast, Moment and BYOL exhibit less consistent prediction distributions, leading to higher error rates.
@@
Skin tone analysis. Understanding the impact of PPG models across skin tones is crucial for practical use (Bent et al., 2020; Sjoding et al., 2020). Therefore, we examine BP estimation performance across skin tones. As shown in Figure 9, PaPaGei-S achieves the best BP estimation across light tones. Across dark tones, we notice that BYOL and REGLE obtain the lowest MAE for Systolic BP and Diastolic BP, however, there are no significant differences across all models. While PaPaGei-S obtains the best overall performance, additional work is necessary to improve robustness on darker skin tones.
6 Conclusion and Future Work

### v2

@@ 4.1 Pre-training
Datasets. We pre-train PaPaGei on three datasets: (1) VitalDB (Lee et al., 2022), which includes PPG signals collected during surgery from the patient’s finger ($f$=500Hz), (2) the MIMIC-III waveform database matched subset (Johnson et al., 2016), where finger-tip PPG data is collected from an ICU monitor ($f$= 125Hz), and (3) the Multi-Ethnic Study of Atherosclerosis (MESA) sleep sub-study (Zhang et al., 2018; Chen et al., 2015), which provides PPG data obtained through finger-tip polysomnography ($f$= 256Hz). In total, we have 13.5K participants with 20M segments (Table 1).
Table 1: PaPaGei’s pre-training datasets.
Dataset #Participants #Segments Hours
@@
Pre-processing. To curate single-channel PPG signals across all datasets, we perform the following steps: (1) Apply a 4th-order Chebyshev bandpass filter with low and high pass cut-offs set at 0.5Hz and 12Hz, respectively (Lapitan et al., 2024; Liang et al., 2018c); (2) Segment the signal into 10-second windows ((Orphanidou, 2018; Koteska et al., 2022) use 10s windows whereas larger studies use 30s (Ding et al., 2024) and 60s Abbaspourazad et al. (2023)); (3) Detect flatline segments and remove any segment where more than 25% of the data is flat (BioBSS Documentation, 2023); (4) Normalize the segments using Z-score (Temko, 2017; Zhou et al., 2017); and (5) Resample the segments to 125Hz (the lowest sampling rate of our pre-training datasets, MIMIC-III).
Table 2: PaPaGei’s evaluation datasets. Gray lines are unseen during training (out-of-domain). For those used for pre-training, we keep a held-out test-sets and use labels. Task Types are: B=binary, R=regression, M-#classes= muticlass classification.
#ID Dataset Task (Task Type) #Subj.(#Samp.)
@@
Implementation. We adopt a ResNet-style CNN encoder, following (Ding et al., 2024). Abbaspourazad et al. (2023) also utilize an EfficientNet-style CNN. Our model has 18 convolutional blocks, starting with a filter size of 32, which doubles every 4 blocks. The projection layer is a single FC layer, generating a 512-d embedding. In the PaPaGei-S variant, the expert block ($M_{1}$ $M_{2}$) uses three parallel FCNNs, each with two FC layers, resulting in a 128-d embedding. For augmentations, PaPaGei-P uses cropping (0.50), negation (0.20), flipping (0.20), and scaling (0.40). PaPaGei-S uses cropping (0.25) and Gaussian noise (0.25). PaPaGei-S avoids augmentations that alter PPG’s morphology. We set $\alpha=0.6$ and train on eight V100 GPUs for 15,000 steps (lr$=10^{-4}$), with PaPaGei-P and PaPaGei-S having 5M and 5.7M parameters, respectively, while previous works use model sizes of 3.3M (Abbaspourazad et al., 2023) (we study scaling in Section 5.2).
@@ 4.2 Downstream Tasks
To evaluate the effectiveness of PaPaGei, we benchmark it against a diverse set of datasets, tasks, and baselines, chosen for their large size and clinical relevance (where applicable) [^2] https://peterhcharlton.github.io/post/ppg_datasets/. A description of the tasks with their corresponding #ID is provided in Table 2, with further details in Appendix §B. As a motivation, identifying patient risk factors is crucial for hospitals to allocate resources effectively. To address this, we evaluate several indicators, including ICU admission (T1), type of operation (T2), mortality (T3), and smoking status (T4). For sleep apnea diagnosis, the American Academy of Sleep Medicine recommends using the Apnea/Hypopnea Index (AHI) with at least 3% or 4% oxygen desaturation as a key metric (Ruehland et al., 2009). Thus, we predict AHI at 3% and 4% desaturation thresholds (T5 & T6) and classify sleep-disordered breathing (T15). For pregnancy outcomes, changes in gestational age and pregnancy stage are linked to risks like hypertensive disorders and small-for-gestational-age delivery (Bouariu et al., 2022; Wu et al., 2020; Crump et al., 2023), enabling us to classify pregnancy stage (T7) and predict gestational age (T8). In cardiovascular health, we estimate systolic (T9 & T11) and diastolic (T10 & T12) blood pressure (BP) using two datasets. While PPG-BP (T11 & T12) provides high-frequency, short PPG signals, the VV dataset helps explore skin tone’s influence on BP estimation. We also assess hypertension classification (T14), average seated heart rate (T13), and continuous heart rate during activities (T19), along with activity classification (T20). In the emotion domain, we classify PPG signals into mood disturbance levels (T16), valence (T17), and arousal (T18).
@@ 4.3 Baselines
We benchmark PaPaGei’s performance against competitive baselines. As open-source foundation models designed for physiological signals, PaPaGei is compared to recent time-series FMs: Chronos (Ansari et al., 2024) and MOMENT (Goswami et al., 2024). To evaluate the merits of our SSL framework, we also compare PaPaGei with common SSL methods (trained from scratch) such as SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), and TF-C (Zhang et al., 2022). In addition, to assess model generalizability on PPG signals, we compare against REGLE, a model pre-trained on UK Biobank’s PPG signals (Yun et al., 2024). As a simple baseline, we employ a random forest trained on statistical features extracted from the PPG signal, including mean, median, maximum, minimum, and the 25th, 50th, and 75th percentiles (”Stat. Features”). This task-specific approach serves as a benchmark for comparison with more advanced techniques.
@@ 4.4 Linear Evaluation
Initially, we split the in-domain and out-of-domain datasets into training, validation, and test sets at 80/10/10 and 60/20/20 ratios. The splitting is performed at the subject level ensuring no overlap between individuals across the sets. The models are evaluated by extracting feature representations from resampled data (125Hz) and applying linear probing for each task. For binary classification tasks, we employ a logistic regression model, with performance measured by the AUROC score. For regression tasks, ridge regression is used, and performance is evaluated based on the mean absolute error (MAE). Regression tasks are aggregated using the symmetric mean absolute percentage error (sMAPE). Multi-class classification tasks are trained using a random forest model, with accuracy as the evaluation metric. To ensure robustness, we compute 95% confidence intervals through bootstrapping (500 sampling runs with replacement). More details are provided in the Appendix §A.
@@ 5.1 Overall Performance
Table 3: Downstream comparison against pre-trained models. Feature extraction parameters are indicated next to each name. 95% CIs are reported in square brackets and the best value is bolded.
REGLE (0.07M) Chronos (200M) Moment (385M) PaPaGei-P (5M) PaPaGei-S (5M)
@@
Average MAE (sMAPE) 12.09 $\pm$ 3.83 (15.23%) 11.60 $\pm$ 3.60 (14.20%) 10.43 $\pm$ 3.46 (13.82%) 10.92 $\pm$ 4.25 (14.09%) 10.12 $\pm$ 3.47 (13.34%)
In general, from Figure 3, we observe that PaPaGei is more accurate across many tasks indicated by the larger AUROC area and smaller MAE area. Table 3 presents a more detailed comparison
@@
between PaPaGei and other pre-trained models. For classification tasks, PaPaGei-S achieves the highest average AUROC of 0.67, outperforming other models across several tasks, particularly in ICU Admission (0.79), Hypertension (0.77), and Sleep Disordered Breathing (0.70). In regression tasks, PaPaGei-S again demonstrates strong performance, achieving the lowest average MAE (10.12), particularly in tasks related to Apnea/Hypopnea Index and BP measurements. REGLE, a small model trained on a large PPG dataset, generally underperforms compared to other models, suggesting its compact size may limit learning complex patterns. Chronos obtains good performance in predicting mortality, pregnancy stage, and smoking, likely due to their slower rate of change and reduced reliance on granular PPG-specific features. General-purpose models suffice for these high-level outcomes. However, tasks requiring finer PPG-specific granularity, such as heart rate prediction, blood pressure estimation, or sleep apnea, benefit from PaPaGei’s specialized feature extraction. Notably, PaPaGei-S consistently outperforms PaPaGei-P, highlighting the advantages of signal morphology objectives in enhancing predictive accuracy.
Table 4 presents a comparison against three SSL methods and a baseline model trained on statistical features. In classification tasks, PaPaGei-S again shows the highest average AUROC, outperforming all others. SimCLR, BYOL, and TF-C generally outperform the statistical feature baseline but fall short of PaPaGei-S’s performance. TF-C shows competitive results in some tasks, achieving the highest AUROC for Mood Disturbance and Valence. For regression tasks, PaPaGei-S again achieves the lowest average MAE. SimCLR, BYOL, and TF-C show mixed results, as each excels in different tasks. SimCLR comes second in estimating Avg HR, while BYOL does so in Systolic BP (VV). The statistical feature baseline generally underperforms compared to the advanced methods across most tasks. PaPaGei-P, while not consistently outperforming PaPaGei-S, shows strong results that are often competitive with or better than other contrastive learning methods. Overall, both PaPaGei variants offer robust performance across a wide range of tasks.
Table 4: Downstream comparison against baseline and SSL methods. Feature extraction parameters are indicated next to each name. 95% CIs are reported in square brackets and the best value is bolded. Implementation details are in Appendix §A.
Stat. Features SimCLR (5M) BYOL (5M) TF-C (10M) PaPaGei-P (5M) PaPaGei-S (5M)
@@ 5.2 Ablation Studies
Pre-training data ablation. We evaluate PaPaGei-S using different pre-training data
Figure 5: Ablation on pre-training data. Average performance across tasks for models trained on: V (VitalDB), M (MESA), and M-III (MIMIC-III). The mean value is displayed above the plots. The Wilcoxon signed rank test is applied to evaluate significance between the All dataset and the rest ($**:p<0.05\;\text{and}\;*:0.05\leq p<0.10$).
combinations. As shown in Figure 5, performance on downstream tasks improves with more upstream data, with the best results achieved when using all three datasets. Notably, MESA outperforms the others despite having the fewest participants but the highest number of segments. This supports findings from language models (Dubey et al., 2024) and wearable sensing research (Narayanswamy et al., 2024), indicating that the volume of segments or hours contributes more to performance than the number of users.
PaPaGei-S component ablation. We assess the impact of PaPaGei-S components. Figure 6 shows that the full model (0.67, 10.12) consistently outperforms individual components in both mean and median metrics. On average, sVRI (0.64, 10.35) outperforms the combinations of sVRI + SQI (0.62, 10.80) and sVRI + IPA (0.64, 10.73). Our results indicate that combining SQI and IPA yields greater benefits compared to their individual contributions.
@@
Figure 6: PaPaGei-S component ablation study (a, b) and scaling analysis (c). (Left) The boxplot shows the performance of PaPaGei-S components across all tasks. The Wilcoxon signed rank test is applied to evaluate pair-wise significance ($**:p<0.05\;\text{and}\;*:0.05\leq p<0.10$). (Right) Heatmap ranks of PaPaGei-S models with 5M, 35M, and 139M parameters (rank 1 denotes the best performance). Detailed results in Table 13.
Model size and scaling analysis. We investigated the impact of model size on performance by training PaPaGei-S-35M and PaPaGei-S-139M with 35M and 139M parameters, respectively.
Figure 7: Pair-wise inter-participant embedding distances for SDB.
Both models share the same number of layers, but the 35M model uses a 32-filter size while the 139M model uses 64. As shown in 6(c), the smallest model (5M parameters) consistently outperformed larger models on all but one task. This suggests the 5M model is better suited for our pre-training datasets, aligning with prior findings on the proportionality between data and model size (Narayanswamy et al., 2024). While the 139M model surpassed the 35M, it still lagged behind the 5M, indicating that wider models may improve performance in classification tasks, likely due to the contrastive learning objective. Nevertheless, our scaling analysis shows a non-monotonic trend, indicating other factors strongly influence performance.
Effect of Demographics. We evaluate the effect of demographics (age, sex) and PPG-specific features (sVRI, IPA, SQI) in Appendix §E. In demographics prediction (Table 16), PaPaGei-S achieves 7.78 MAE in age regression, 0.85 accuracy in age classification, and 0.79 accuracy in sex classification. While our results trail larger closed studies (Abbaspourazad et al., 2023) by 2.18, 0.05, and 0.13 for segment-level SSL, and by 5.59, 0.12, and 0.25 for patient-level SSL, they mark an advancement in open-source efforts. These findings indicate that patient-level positive pair selection in SSL better captures demographic-related features for downstream prediction. Moreover, the reduced performance of PaPaGei-S can be attributed to evaluations conducted on diverse device setups, as opposed to a single device configuration. Our ablation study (Table 15) shows that PaPaGei-S outperforms the demo + PPG in 14 out of 18 tasks, particularly in tasks with real-time dependence such as heart rate estimation. Importantly, including demographics in addition to PaPaGei-S creates a stronger model. These findings emphasize that demographic features complement rather than compete with PaPaGei-S, showcasing the potential of integrating PaPaGei’s advanced feature extraction capabilities with demographic context to improve task outcomes.
@@ 5.3 Case Studies
Inter-participant embeddings. Figure 7 shows the distribution of pair-wise embedding distances across participants in the SDB dataset (Kiyasseh et al., 2021). SimCLR and BYOL exhibit sharper peaks at lower distances, indicating that participants are more closely clustered within the embedding space. This could be interpreted as a mild form of mode collapse, where the model does not fully capture the individual differences between participants. TF-C demonstrates a more balanced distribution, with both large and small peaks, suggesting it captures both similarities and some variation between participants. In contrast, PaPaGei-S provides the widest dispersion of embeddings, highlighting its ability to capture a broader range of features that may be valuable for distinguishing between participants’ medical conditions.
Figure 8: Regression plots and prediction distribution of different models compared to ground truth for AHI $>$ 3%. $R^{2}$ is the coefficient of determination and $m$ is the correlation slope.
Regression predictions. From Figure 8, compared to the pre-trained and SSL baseline, PaPaGei-S demonstrates steeper slopes ($m$) and higher $R^{2}$ values, reflecting a stronger alignment between predictions and true values. Additionally, the prediction distribution for AHI indicates that SimCLR and Chronos tend to regress more toward the mean, while PaPaGei-S achieves a wider distribution base, highlighting its capacity to capture left tail better. Additional plots are shown in Appendix §F.
@@
Skin tone analysis. We examine BP estimation performance across skin tones because it is crucial for practical use (Bent et al., 2020). As shown in Figure 9 (More details in Figure 26), PaPaGei-S achieves the best BP estimation across light tones. Across dark tones, we notice that BYOL and REGLE obtain the lowest MAE for Systolic BP and Diastolic BP. However, identifying a single model that performs best across all skin tones remains challenging. While PaPaGei-S obtains the best overall performance, additional work is necessary to improve robustness on darker skin tones.
Figure 9: Skin tone analysis for Blood Pressure estimation (VV dataset)
6 Discussion & Conclusion